# Rehearsal-Free Continual Federated Learning with Synergistic Regularization

### v1

@@ 5.2 Performance Overview
Table 4: Evaluation of various methods in terms of the communication rounds to reach the best test accuracy. We report the sum of communication rounds required to achieve the best performance on each task and evaluate with the trade-offs between communication rounds and accuracy. We denote "$\Delta$" as the difference between the accuracy improvement percentage and the round increase percentage of FedSSI and other baselines.
@@
Method Time Cost Main Cause
FedAvg 3.13h -
FL+LwF 4.46h Knowledge Distillation
FL+EWC 5.75h Fish Information Matrix
FL+OGD 7.09h Orthogonal Projection
FedSSI 3.75h Surrogate Loss
@@
Data Heterogeneity. Figure 3 displays the test accuracy with different levels of data heterogeneity on four datasets. As shown in this figure, all methods achieve an improvement in test accuracy with the decline in data heterogeneity, and FedSSI consistently achieves a leading and stable improvement in performance with different levels of data heterogeneity.
Resource Consumption. Table 5 lists the training time of the four methods mentioned above. Although most traditional regularization algorithms are not suitable for resource-constrained edge devices in the CFL scenario due to their computational efficiency overhead, we also observe that a few methods that could potentially be deployed in CFL, such as SI, have a relatively low time overhead, achieving significant performance improvements with less than 20% increase in training time compared to fine-tuning (FedAvg).
Table 4 compares the communication efficiency of various methods by measuring the trade-offs between communication rounds and accuracy. Compared to other CFL methods, FedSSI may introduce additional communication rounds but can achieve better performance with a cost-effective trade with a large $\Delta$ value. We observed that FedSSI can achieve a significant $\Delta$ value for all datasets except Digit10, indicating that our communication method is efficient and can bring considerable performance improvements. For Digit10, a possible reason is that the dataset itself has simple features (one-channel images), and all methods can achieve relatively good accuracy. In such cases, the percentage increase in accuracy by FedSSI is relatively small. However, achieving a high-performance improvement on a simple dataset is not easy in itself. In the next version, we will consider using a smaller network model (as learning Digit10 with a CNN is sufficient) for verification. The communication rounds for each task of different datasets are listed in Table 5.1.
@@
Hyper-parameter. Then, we conduct more research on the setting of hyper-parameter $\lambda$. In our framework, we modify the $\lambda$ value to adjust the global and local information proportion in PSM. As shown in Table 3, we select three different $\lambda$ values with four different data heterogeneity settings and evaluate the final test accuracy on three datasets. Experimental results show that the value of $\lambda$ should be chosen accordingly under different data heterogeneity. Nevertheless, results exhibit the same trend: as the degree of data heterogeneity increases, FedSSI performs better while $\lambda$ decreases as PSM contains more global information. Empirically, striking a balance between global and local information is the key to addressing the data heterogeneity in CFL. Vice versa, as $\alpha$ increases, the data distribution on the client side becomes more IID. At this point, the clients require less global information and can rely more on their local information for caching important samples. Although we cannot directly relate $\alpha$ and $\lambda$ with a simple formula due to the complexity of the problem, even in specialized research on personalized federated learning (PFL), methods such as Gaussian mixture modeling are relied upon. we can empirically and theoretically judge there exists a positive correlation between $\alpha$ and $\lambda$ with Proposition 1 and Table 3.
Quantitative Analysis. Figure 4 shows the qualitative analysis of the number of incremental tasks $n$ on three class-incremental datasets. According to these curves, we can easily observe that our model performs better than other baselines across all tasks, with varying numbers of incremental tasks. It demonstrates that FedSSI enables clients to learn new incremental classes better than other methods.
6 Conclusion and Future Work
@@
In this paper, we introduced FedSSI, a novel regularization-based method for continual federated learning (CFL) that mitigates catastrophic forgetting and manages data heterogeneity without relying on data rehearsal or excessive computational resources. Our extensive experiments show that FedSSI outperforms existing CFL approaches, achieving superior accuracy and stability in resource-constrained environments. This work paves the way for more practical and efficient CFL deployments in real-world scenarios.
Although we have invested in the effectiveness of regularization-based methods over the CFL scenarios without relying on data rehearsal or excessive computational resources, the overhead of other training resources should be taken into account. To deploy the FL system in practical settings, it is necessary to consider resource factors such as training efficiency, model capacity, and even sparsely labeled data. In the future, we seek to work a step forward in this field.
Appendix
@@ Appendix C Configurations
Appendix D Limitation
The limitation of FedSSI is that the PSM may introduce additional storage. However, the model for an FL task is practically not large as edge clients are mostly resource-limited, and the memory demand by PSM is relatively similar to existing methods like FedProx as to manipulate over an extra model that has the same size as the federated model. In addition, we propose a possible strategy to address this issue: If an FL task has large models and limited memory left on the edge, a prerequisite can be reasonably assumed satisfied: transmission capacity (probably after optimization) is sufficient for the system. In such a case, we can record the PSM model in the server, which is only downloaded and updated locally to compute parameter contributions and is then uploaded to the server again.
Appendix E Additional Results
In this section, we first provide the experiments about scalability and bandwidth constraints to validate the effectiveness of our method. Then, we provide more details about the experiment results of each task. We record the test accuracy of the global model at the training stage of each task and the communication rounds required to achieve the corresponding performance.
E.1 Detailed Results of Scalability and Bandwidth Constraints
@@
FedAvg 36.68 59.17 68.12 80.34 29.02 53.89 63.06 78.27
FL+EWC 38.31 60.19 69.25 81.54 28.86 54.52 62.92 78.80
Re-Fed 38.08 59.02 67.85 79.85 28.82 52.84 61.53 76.69
FOT 40.18 61.41 68.54 79.70 31.31 55.56 62.41 76.32
FedSSI 42.58 62.65 72.09 82.49 32.95 56.12 64.91 79.10
E.2 Detailed Results of Test Accuracy
@@
E.3 Detailed Results of Communication Round.

### v2

@@ 5.1 Experiment Setup
Table 4: Evaluation of various methods in terms of the communication rounds to reach the best test accuracy. We report the sum of communication rounds required to achieve the best performance on each task and evaluate with the trade-offs between communication rounds and accuracy. We denote ”$\Delta$” as the difference between the accuracy improvement percentage and the round increase percentage of FedSSI and other baselines.
@@
5.2 Performance Overview
Test Accuracy. Table 2 shows the test accuracy of various methods with data heterogeneity across six datasets. We report both the final accuracy and average accuracy of the global model when all clients finish their training on all tasks. FOT outperforms other baselines on three class-IL datasets as the orthogonal projection in FOT exhibits effective training with obvious boundaries between streaming tasks. However, its performance experiences a significant decline with domain-IL datasets, where the classes of each task remain unchanged. All regularization-based methods work and CFL models with data rehearsal fail to achieve ideal results with limited memory buffer. FedSSI achieves the best performance in all cases by up to 11.52% in final accuracy. More discussions and results on model performance and communication efficiency are available in Appendix D.
@@ 5.2 Performance Overview
Data Heterogeneity. Figure 3 displays the test accuracy with different levels of data heterogeneity on four datasets. As shown in this figure, all methods improve test accuracy with the decline in data heterogeneity, and FedSSI consistently achieves a leading and stable improvement in performance with different levels of data heterogeneity.
@@
Hyper-parameter. Then, we conduct more research on the setting of hyper-parameter $\lambda$. In our framework, we modify the $\lambda$ value to adjust the global and local information proportion in PSM. As shown in Table 3, we select three different $\lambda$ values with four different data heterogeneity settings and evaluate the final test accuracy on three datasets. Experimental results show that the value of $\lambda$ should be chosen accordingly under different data heterogeneity. Nevertheless, results exhibit the same trend: as the degree of data heterogeneity increases, FedSSI performs better while $\lambda$ decreases as PSM contains more global information. Empirically, striking a balance between global and local information is the key to addressing the data heterogeneity in CFL. Vice versa, as $\alpha$ increases, the data distribution on the client side becomes more IID. At this point, the clients require less global information and can rely more on their local information for caching important samples. Although we cannot directly relate $\alpha$ and $\lambda$ with a simple formula due to the complexity of the problem, even in specialized research on PFL, methods such as Gaussian mixture modeling are relied upon. we can empirically and theoretically judge there exists a positive correlation between $\alpha$ and $\lambda$ with Proposition 1 and Table 3.
Resource Consumption Table 4 compares the communication efficiency of various methods by measuring the trade-offs between communication rounds and accuracy. Compared to other CFL methods, FedSSI may introduce additional communication rounds but can achieve better performance with a cost-effective trade with a large $\Delta$ value. We observed that FedSSI can achieve a significant $\Delta$ value for all datasets except Digit10, indicating that our communication method is efficient and can bring considerable performance improvements. For Digit10, a possible reason is that the dataset itself has simple features (one-channel images), and all methods can achieve relatively good accuracy. In such cases, the percentage increase in accuracy by FedSSI is relatively small. However, achieving a high-performance improvement on a simple dataset is not easy in itself. In the next version, we will consider using a smaller network model (as learning Digit10 with a CNN is sufficient) for verification. The communication rounds for each task of different datasets are listed in Section 5.1.
6 Conclusion
@@
In this paper, we introduced FedSSI, a novel regularization-based method for continual federated learning (CFL) that mitigates catastrophic forgetting and manages data heterogeneity without relying on data rehearsal or excessive computational resources. Our extensive experiments show that FedSSI outperforms existing CFL approaches with superior accuracy and stability in resource-constrained environments.
Acknowledgments
This work is supported by the National Key Research and Development Program of China under grant 2024YFC3307900; the National Natural Science Foundation of China under grants 62376103, 62302184, 62436003 and 62206102; Major Science and Technology Project of Hubei Province under grant 2024BAA008; Hubei Science and Technology Talent Service Project under grant 2024DJC078; Ant Group through CCF-Ant Research Fund; and Fundamental Research Funds for the Central Universities under grant YCJJ20252319. The computation is completed in the HPC Platform of Huazhong University of Science and Technology.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Configurations
Appendix D Additional Results
In this section, we first provide the experiments about training time cost and quantitative analysis. Then, we conduct experiments on scalability and bandwidth constraints to validate the effectiveness of our method. Finally, we provide more details about the experiment results of each task. We record the test accuracy of the global model at the training stage of each task and the communication rounds required to achieve the corresponding performance.
D.1 Detailed Results of Time Cost and Quantitative Analysis
Table 6 lists the training time of the four methods mentioned above. Although most traditional regularization algorithms are not suitable for resource-constrained edge devices in the CFL scenario due to their computational efficiency overhead, we also observe that a few methods that could potentially be deployed in CFL, such as SI, have a relatively low time overhead, achieving significant performance improvements with less than 20% increase in training time compared to fine-tuning (FedAvg).
Figure 4 shows the qualitative analysis of the number of incremental tasks $n$ on three class-incremental datasets. According to these curves, we can easily observe that our model performs better than other baselines across all tasks, with varying numbers of incremental tasks. It demonstrates that FedSSI enables clients to learn new incremental classes better than other methods.
Method Time Cost Main Cause
FedAvg 3.13h -
FL+LwF 4.46h Knowledge Distillation
FL+EWC 5.75h Fish Information Matrix
FL+OGD 7.09h Orthogonal Projection
FedSSI 3.75h Surrogate Loss
D.2 Detailed Results of Scalability and Bandwidth Constraints
@@
FedAvg 18.67 45.80 55.91 70.37 29.02 53.89 63.06 78.27
FL+EWC 19.93 46.38 56.82 70.40 28.86 54.52 62.92 78.80
Re-Fed 19.44 44.08 54.91 66.24 28.82 52.84 61.53 76.69
FOT 21.26 47.02 56.06 69.69 31.31 55.56 62.41 76.32
FedSSI 23.61 47.14 59.35 71.27 32.95 56.12 64.91 79.10
D.3 Detailed Results of Test Accuracy
@@
D.4 Detailed Results of Communication Round.